The number of bacteria N(t)present in a culture at time t hours is given by the following.
N(t) = 1600(2)^t
Find the number of bacteria for the following values of t.
(a) t = 0 hours

bacteria
(b) t = 6 hours

bacteria

Answer by fractalier(6550)

For the formula
N(t) = 1600(2)^t
we have
N(0) = 1600(2)^0 = 1600 bacteria
N(6) = 1600(2)^6 = 102400 bacteria